Gilbert popped in, but toddled off to Winchcombe to help Fred on the siphon restoration because there is little to do on 2807.
One new issue has been logged. Here is the list of outstanding issues and their status:
1 J cocks are stuck [A last-year’s issue: Probably in the cab; in ‘open’ position]
2 LH front & rear top mud hole doors & right-hand top blowing [Marked as “Monitor”]
7 Safety valves lift at 215 lbs [We will verify and adjust at next wash-out]
12 RH piston gland packing blowing [This is not a significant leak - Monitor]
18 L/H clack valve/injector steam valve blowing slightly [We’ll take a look when she’s next in steam].
Chaps from the Loco Dept, supervised by Brian, cleaned out the grate and the smokebox. They also cleaned the bottom half of the loco including the motion.
Brian cleaned up the water scoop column because the fracture in the casing became visible again. Bruce re-painted the column.
John G wire-brushed the queue of rail chairs in the boot scraper production line, and Bruce painted their bottoms.
I finished off a load of boot scrapers - painted the lettering on them and fitted the brushes.
Bruce & Brian joined in as Dinmore Manor attempted to weigh their new tender. The usual amusing discrepancies arose: how can it be heavier one side than the other; how come the weights change when you run it up & down the track and weigh it again. 😊
2807 is next in service at the gala (27th - 29th May). Today was ‘Lego day’, and with parking on the field plus nothing much to do on 2807, we had the day off.
Well, I did some painting of chairs and cut up a multitude of wooden wedges for the boot scrapers (to hold the brushes in place). Pete and Alex from the Loco dept cleaned 2807’s grate. Pete reported that there appear to be a number of stays weeping; ditto at the corner patches, and one tube could be partially blocked.
Meanwhile, I had done some investigation of the historic report of “J cocks stuck”. I found out which ones … and, yes I couldn’t close them.